Can you tell me how to do this



5. Fill in the chart given this linear equation:
3x + y =10
x y=
-2
-1
0
1
2

Answer by ilana(305)

First, write the equation in terms of y. y = 10-3x, or y = -3x + 10.
Now substitute each x. When x=-2, y = 10 - 3*(-2), so y = 10 - (-6), or y=10+6, so y=16. Substitute each value of x to complete the chart.